OpenStudy (anonymous):

0.5 divided by 1/2 isnt it 0.25???

OpenStudy (anonymous):

no its 1

OpenStudy (anonymous):

Nope 1/2 /(1/2) = 1/2 * (2/1) = 1

OpenStudy (anonymous):

0.5 times 1/2 = .25

OpenStudy (amistre64):

1/2//1/2 = 1 ... anything over itself = 1 lol

OpenStudy (anonymous):

oh okay thanks

OpenStudy (anonymous):

((93 x 26)divide by 2) x 0 +19

OpenStudy (amistre64):

19 then

OpenStudy (amistre64):

(...yada..)0 + 19 = 19